it says to find x and the measure of each side of the triangle. triangle DEF is isosceles, angle D is the vertex angle, DE=x+7, Df=3x-1, and EF=2x+5
Answer by mamiya(56)

it says to find x and the measure of each side of the triangle. triangle DEF is isosceles, angle D is the vertex angle, DE=x+7, Df=3x-1, and EF=2x+5
D is a vertex angle, so DE=DF
thus, x+7=3x-1
8=2x
x=4
DE= x+7
=4+7
=11
DF= 3X-1
=3*4 -1
=11
EF= 2x-5
= 2*4 - 5
= 3